OpenStudy (timinthor):

What is the standard for of this equation? 3y – 7 = –5x

OpenStudy (igreen):

Standard form is: \(\sf y = mx + b\) So first, add 7 to both sides, can you do that?

OpenStudy (timinthor):

Yes one sec

OpenStudy (timinthor):

So would you have 3y=-5x+7

OpenStudy (dayakar):

standard form any linear equation in two variables x and y is ax+by+c=0

OpenStudy (timinthor):

So would the answer be C

OpenStudy (timinthor):

@iGreen, So the answer is C?

OpenStudy (igreen):

You didn't give options.

OpenStudy (igreen):

Now we have to divide 3 to both sides.

OpenStudy (igreen):

Oh, that's right..I got mixed up..standard form is \(\sf Ax + By = C\)

OpenStudy (timinthor):

oh wait sorry my bad A. 5x + 3y = 7 B. 3y – 7 = –5x C. 3y = –5x + 7 D. 5x + 3y – 7 = 0

OpenStudy (igreen):

So we actually have to add 5x to both sides now.

OpenStudy (igreen):

\(\sf 3y = -5x + 7\) Add 5x to both sides.

OpenStudy (timinthor):

So then 3y+5x=7 ?

OpenStudy (timinthor):

Which would make it A, because 3y+5x is the same as 5x+3y right?

OpenStudy (igreen):

Yep, you got it!

OpenStudy (timinthor):

Thank you!